OpenStudy (anonymous):

Solve -4x - 5 > -9 and describe the graph of the solution?

OpenStudy (amistre64):

since x=0 makes a true statment; id say darken the section with the 0 on it

OpenStudy (amistre64):

to find the starting point; just +5 and divide off the -4 the division of a negative makes us flip the sign tho

OpenStudy (amistre64):

which really doesnt matter since the 0 part is the answer either way :)

OpenStudy (amistre64):

-9+5 = -4 -4/-4 = 1 circle the 1 and go left

OpenStudy (anonymous):

Would that be an open or closed circle?

OpenStudy (amistre64):

open since its the ">" symbol. no solid line beneath it means no solid point